# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

### v2

@@ Appendix E Full results
space_ga 3% 0% 3% 0.92 2.15 0.23 -19% -24% 0.51
space_ga 8% -1% 9% 0.91 2.04 0.22 -15% -18% 0.54
space_ga 10% -1% 10% 0.91 2.04 0.22 -16% -17% 0.54
space_ga 11% -2% 11% 0.91 2.01 0.22 -14% -17% 0.55
space_ga 12% -3% 13% 0.90 1.96 0.21 -14% -15% 0.55
cpu_activity 24% -8% 26% 0.91 0.48 0.09 -15% -22% 0.98
cpu_activity 24% -8% 25% 0.91 0.49 0.09 -17% -22% 0.97
cpu_activity 30% -9% 30% 0.90 0.45 0.09 -13% -18% 0.98
cpu_activity 29% -8% 28% 0.90 0.47 0.09 -13% -18% 0.98
cpu_activity 27% -8% 27% 0.91 0.48 0.09 -14% -20% 0.98
naval_propulsion_plant 265% -9% 75% 0.91 0.58 0.17 -268% -371% 0.97
naval_propulsion_plant 205% -9% 70% 0.90 0.69 0.20 -332% -419% 0.96
naval_propulsion_plant 178% -9% 67% 0.91 0.74 0.22 -382% -464% 0.95
naval_propulsion_plant 198% -9% 69% 0.91 0.69 0.20 -336% -422% 0.96
naval_propulsion_plant 248% -10% 74% 0.90 0.58 0.17 -283% -368% 0.97
miami_housing 70% -12% 48% 0.86 0.66 0.08 -12% -22% 0.90
miami_housing 63% -12% 46% 0.86 0.68 0.08 -9% -21% 0.90
miami_housing 68% -13% 48% 0.86 0.65 0.08 -9% -20% 0.90
miami_housing 71% -13% 49% 0.86 0.63 0.08 -7% -18% 0.90
miami_housing 72% -12% 49% 0.86 0.65 0.08 -12% -21% 0.89
kin8nm 12% -9% 19% 0.88 2.20 0.43 -32% -36% 0.50
kin8nm 9% -9% 17% 0.88 2.28 0.45 -36% -40% 0.46
kin8nm 11% -10% 18% 0.88 2.24 0.44 -34% -37% 0.49
kin8nm 9% -9% 16% 0.89 2.29 0.45 -34% -36% 0.49
kin8nm 7% -9% 15% 0.88 2.33 0.46 -37% -40% 0.47
concrete_compressive_strength 25% -15% 32% 0.79 1.02 0.21 -59% -85% 0.78
concrete_compressive_strength 9% -12% 19% 0.81 1.19 0.24 -71% -96% 0.76
concrete_compressive_strength 4% -11% 14% 0.83 1.22 0.25 -62% -92% 0.77
concrete_compressive_strength -2% -9% 8% 0.83 1.31 0.27 -77% -110% 0.74
concrete_compressive_strength 11% -12% 20% 0.81 1.14 0.23 -64% -88% 0.77
cars 3% 15% -11% 0.83 0.54 0.09 3% 2% 0.95
cars 16% 18% -2% 0.85 0.52 0.09 3% 0% 0.95
cars 21% 13% 7% 0.80 0.47 0.08 4% -1% 0.95
cars 6% 17% -10% 0.84 0.56 0.10 -2% 0% 0.94
cars 10% 17% -7% 0.85 0.54 0.09 8% 5% 0.95
energy_efficiency -3% -3% 0% 0.63 0.07 0.02 3% 0% 1.00
energy_efficiency -3% 1% -5% 0.66 0.08 0.02 7% 3% 1.00
energy_efficiency -14% 9% -29% 0.70 0.09 0.03 0% 0% 1.00
energy_efficiency 2% -11% 14% 0.59 0.07 0.02 -3% -2% 1.00
energy_efficiency -5% 1% -5% 0.65 0.08 0.02 6% 0% 1.00
california_housing 33% -7% 30% 0.91 1.71 0.41 -14% -19% 0.72
california_housing 28% -7% 27% 0.91 1.74 0.41 -16% -22% 0.70
california_housing 34% -7% 31% 0.90 1.67 0.40 -14% -20% 0.72
california_housing 29% -7% 28% 0.91 1.72 0.41 -16% -22% 0.71
california_housing 31% -8% 30% 0.90 1.69 0.40 -15% -20% 0.71
airfoil_self_noise 8% -21% 27% 0.77 1.20 0.26 -168% -189% 0.74
airfoil_self_noise -16% -14% -3% 0.84 1.68 0.36 -170% -214% 0.67
airfoil_self_noise 20% -8% 24% 0.89 1.22 0.26 -95% -105% 0.84
airfoil_self_noise 62% -18% 49% 0.81 0.82 0.18 -57% -81% 0.89
airfoil_self_noise 18% -13% 26% 0.86 1.19 0.25 -93% -118% 0.85
QSAR_fish_toxicity 11% -10% 19% 0.81 1.82 0.28 -6% -13% 0.53
QSAR_fish_toxicity 13% -10% 21% 0.82 1.79 0.28 -5% -11% 0.54
QSAR_fish_toxicity 14% -10% 21% 0.82 1.82 0.28 -8% -12% 0.51
QSAR_fish_toxicity 11% -8% 17% 0.82 1.79 0.28 -5% -11% 0.53
QSAR_fish_toxicity 9% -8% 16% 0.84 1.84 0.29 -6% -10% 0.52
@@
space_ga 7% -4% 10% 0.79 1.22 0.13 -1% -1% 0.75
space_ga 8% 2% 5% 0.82 1.25 0.14 0% 0% 0.76
space_ga 7% 4% 3% 0.83 1.27 0.14 -3% -3% 0.74
space_ga 8% -1% 8% 0.79 1.24 0.14 0% -1% 0.74
space_ga 9% -2% 10% 0.81 1.27 0.14 1% 0% 0.76
cpu_activity 8% -11% 16% 0.72 0.25 0.05 6% 5% 0.98
cpu_activity 14% -15% 25% 0.68 0.23 0.04 8% 5% 0.98
cpu_activity 8% -8% 15% 0.73 0.24 0.04 8% 7% 0.98
cpu_activity 5% -3% 8% 0.74 0.25 0.05 6% 5% 0.98
cpu_activity 6% -8% 12% 0.70 0.24 0.04 4% 2% 0.98
naval_propulsion_plant 50% 0% 34% 0.92 0.19 0.06 -44% -48% 1.00
naval_propulsion_plant 20% 0% 17% 0.93 0.20 0.06 -35% -13% 1.00
naval_propulsion_plant 3% 1% 3% 0.92 0.26 0.08 -51% -45% 0.99
naval_propulsion_plant 3% 3% 0% 0.94 0.23 0.07 -86% -63% 1.00
naval_propulsion_plant 29% -6% 27% 0.91 0.23 0.07 -9% -11% 1.00
miami_housing 13% -2% 13% 0.78 0.37 0.05 -2% 7% 0.91
miami_housing 10% 3% 6% 0.81 0.40 0.05 -6% 1% 0.91
miami_housing 14% 1% 10% 0.82 0.42 0.05 -2% 3% 0.92
miami_housing 8% 1% 7% 0.82 0.43 0.05 -1% 4% 0.91
miami_housing 3% 6% -4% 0.83 0.44 0.06 -4% 6% 0.91
kin8nm -5% 9% -15% 0.81 0.69 0.13 6% 5% 0.93
kin8nm 15% 1% 12% 0.78 0.62 0.12 7% 7% 0.93
kin8nm 2% 4% -2% 0.82 0.67 0.13 10% 10% 0.94
kin8nm 2% 14% -12% 0.83 0.69 0.14 10% 6% 0.94
kin8nm -3% 12% -15% 0.83 0.71 0.14 4% 5% 0.93
concrete_compressive_strength 89% -14% 54% 0.62 0.30 0.06 4% 10% 0.90
concrete_compressive_strength 82% -21% 57% 0.55 0.29 0.06 9% 15% 0.91
concrete_compressive_strength 120% -17% 62% 0.60 0.28 0.06 18% 19% 0.92
concrete_compressive_strength 100% -33% 66% 0.52 0.27 0.06 9% 12% 0.91
concrete_compressive_strength 101% -30% 65% 0.52 0.28 0.06 18% 20% 0.91
cars 10% 3% 7% 0.77 0.38 0.06 -4% 5% 0.95
cars -23% 28% -65% 0.88 0.54 0.09 1% 1% 0.96
cars 16% -4% 17% 0.65 0.29 0.05 -6% -2% 0.95
cars 2% -6% 8% 0.68 0.34 0.06 -2% -1% 0.95
cars -5% 8% -13% 0.81 0.45 0.08 -5% -2% 0.95
energy_efficiency 93% 59% 17% 0.70 0.07 0.02 37% 38% 1.00
energy_efficiency 83% 27% 29% 0.61 0.06 0.02 28% 32% 1.00
energy_efficiency 109% 28% 39% 0.64 0.06 0.02 35% 32% 1.00
energy_efficiency 65% 26% 24% 0.58 0.07 0.02 31% 31% 1.00
energy_efficiency 60% -15% 47% 0.58 0.06 0.02 31% 32% 1.00
california_housing -13% 9% -25% 0.90 1.33 0.32 7% 8% 0.81
california_housing -15% 6% -24% 0.89 1.27 0.30 7% 9% 0.82
california_housing -11% 9% -23% 0.90 1.29 0.31 8% 11% 0.82
california_housing -14% 8% -26% 0.89 1.31 0.31 7% 8% 0.81
california_housing -18% 9% -33% 0.90 1.32 0.32 7% 9% 0.82
airfoil_self_noise 140% -10% 62% 0.74 0.34 0.07 18% 18% 0.97
airfoil_self_noise 42% 9% 23% 0.73 0.34 0.07 11% 9% 0.97
airfoil_self_noise 43% -1% 31% 0.77 0.36 0.08 22% 22% 0.97
airfoil_self_noise 39% 3% 26% 0.73 0.35 0.07 17% 12% 0.97
airfoil_self_noise 105% -2% 52% 0.77 0.38 0.08 21% 19% 0.97
QSAR_fish_toxicity -6% 8% -14% 0.72 1.40 0.22 -4% 6% 0.52
QSAR_fish_toxicity 0% 17% -17% 0.82 1.50 0.23 6% 11% 0.54
QSAR_fish_toxicity -5% 8% -13% 0.77 1.57 0.24 0% -1% 0.57
QSAR_fish_toxicity -8% 13% -23% 0.78 1.51 0.23 10% 9% 0.61
QSAR_fish_toxicity -12% 10% -25% 0.73 1.49 0.23 5% 11% 0.51
@@
space_ga -12% 5% -20% 0.87 2.25 0.24 -11% -13% 0.35
space_ga -9% 6% -18% 0.87 2.15 0.23 -12% -11% 0.44
space_ga 3% 6% -5% 0.85 1.90 0.21 -1% -6% 0.46
space_ga -8% 4% -14% 0.89 2.06 0.22 -5% -11% 0.54
space_ga 0% 1% -4% 0.82 1.88 0.20 -6% -6% 0.45
cpu_activity 2% 2% -1% 0.79 0.58 0.11 -30% -30% 0.73
cpu_activity 9% 7% -12% 0.87 0.61 0.12 -7% -23% 0.94
cpu_activity -41% 27% -127% 0.96 1.17 0.22 -1% -26% 0.81
cpu_activity -8% 14% -34% 0.94 0.78 0.15 36% 5% 0.88
cpu_activity 12% 8% -12% 0.91 0.66 0.12 -2% -11% 0.81
naval_propulsion_plant 82% -9% 47% 0.76 0.64 0.19 -30% -42% 0.83
naval_propulsion_plant 121% -13% 59% 0.76 0.65 0.19 5% -6% 0.84
naval_propulsion_plant 127% -13% 58% 0.73 0.66 0.20 6% 6% 0.76
naval_propulsion_plant 134% -21% 63% 0.64 0.52 0.15 -7% -11% 0.70
naval_propulsion_plant 58% -25% 53% 0.62 0.64 0.19 -45% -90% 0.73
miami_housing 16% -6% 13% 0.84 0.86 0.11 -4% -4% 0.72
miami_housing -9% 4% -22% 0.90 0.94 0.12 -14% -20% 0.79
miami_housing 25% -11% 17% 0.72 0.60 0.07 -35% -35% 0.54
miami_housing 23% -4% 21% 0.85 0.83 0.10 -7% -12% 0.80
miami_housing 25% -9% 21% 0.78 0.63 0.08 -10% -19% 0.76
kin8nm -3% 2% -12% 0.88 2.41 0.47 0% 1% 0.40
kin8nm -6% 3% -14% 0.89 2.47 0.48 -6% -8% 0.39
kin8nm -4% 2% -10% 0.88 2.48 0.48 -5% -6% 0.37
kin8nm -9% 7% -23% 0.90 2.57 0.50 -5% -7% 0.44
kin8nm -3% 3% -10% 0.87 2.38 0.47 1% 2% 0.41
concrete_compressive_strength -2% 0% -6% 0.71 1.51 0.30 -19% -22% 0.48
concrete_compressive_strength 2% 11% -15% 0.81 1.44 0.29 -4% -3% 0.69
concrete_compressive_strength 4% 11% -12% 0.76 1.44 0.29 -15% -24% 0.50
concrete_compressive_strength 1% 8% -9% 0.77 1.47 0.30 -4% -9% 0.64
concrete_compressive_strength 7% 8% -7% 0.80 1.49 0.30 -37% -44% 0.57
cars -3% 24% -34% 0.85 1.34 0.23 -9% -16% 0.75
cars -35% 9% -98% 0.82 1.78 0.30 -68% -80% 0.55
cars -6% 29% -52% 0.88 1.54 0.26 -8% -14% 0.76
cars -29% 32% -90% 0.81 1.73 0.29 -28% -39% 0.40
cars -5% 34% -40% 0.92 1.49 0.25 -37% -25% 0.81
energy_efficiency 20% 15% -13% 0.93 0.71 0.21 -37% -39% 0.94
energy_efficiency 35% 3% 23% 0.88 0.58 0.17 -15% -24% 0.97
energy_efficiency 37% 7% 19% 0.92 0.63 0.18 -16% -44% 0.96
energy_efficiency -21% 18% -81% 0.95 1.01 0.29 -155% -137% 0.93
energy_efficiency -22% 14% -66% 0.91 0.92 0.27 -705% -520% 0.76
california_housing 30% -8% 28% 0.82 1.52 0.36 -4% -7% 0.61
california_housing 26% -2% 21% 0.89 1.61 0.39 0% 1% 0.73
california_housing 34% -1% 26% 0.90 1.66 0.40 -11% -13% 0.73
california_housing 13% 0% 11% 0.90 1.85 0.44 -3% -3% 0.65
california_housing 27% 0% 21% 0.91 1.70 0.41 -1% -3% 0.73
airfoil_self_noise -8% 4% -14% 0.77 1.70 0.36 -14% -8% 0.52
airfoil_self_noise -18% -2% -20% 0.80 1.77 0.38 -94% -110% 0.50
airfoil_self_noise -12% -7% -11% 0.70 1.86 0.40 -43% -48% 0.19
airfoil_self_noise 34% -5% 29% 0.69 1.02 0.22 -19% -29% 0.38
airfoil_self_noise -6% -6% -5% 0.71 1.58 0.34 -14% -22% 0.43
QSAR_fish_toxicity -3% -4% -1% 0.68 1.56 0.24 -4% -5% 0.38
QSAR_fish_toxicity -1% -4% 2% 0.75 1.83 0.28 -9% -3% 0.35
QSAR_fish_toxicity 3% 9% -8% 0.75 1.66 0.26 4% 4% 0.43
QSAR_fish_toxicity -6% 1% -19% 0.72 1.61 0.25 -7% -3% 0.35
QSAR_fish_toxicity -1% -3% -2% 0.74 1.84 0.29 -5% -3% 0.30

### v3

@@ Appendix E Full conformalized results
space_ga -8% 2% -11% 0.95 2.59 0.28 -13% -15% 0.55
space_ga -3% 0% -3% 0.95 2.77 0.30 -14% -17% 0.55
space_ga -2% 1% -4% 0.96 2.61 0.28 -13% -13% 0.57
space_ga -9% 1% -11% 0.96 2.81 0.30 -15% -18% 0.54
space_ga -7% 0% -8% 0.96 2.79 0.30 -19% -24% 0.51
cpu_activity 11% -4% 13% 0.95 0.57 0.11 -13% -20% 0.98
cpu_activity 14% -5% 16% 0.94 0.54 0.10 -13% -18% 0.98
cpu_activity 9% -4% 11% 0.95 0.58 0.11 -13% -18% 0.98
cpu_activity 10% -4% 13% 0.94 0.57 0.11 -15% -22% 0.97
cpu_activity 10% -4% 13% 0.94 0.57 0.11 -16% -22% 0.97
naval_propulsion_plant 131% 1% 56% 0.96 0.96 0.28 -320% -418% 0.96
naval_propulsion_plant 144% 1% 59% 0.96 0.88 0.26 -286% -369% 0.97
naval_propulsion_plant 174% 0% 63% 0.96 0.81 0.24 -257% -348% 0.97
naval_propulsion_plant 127% 0% 56% 0.96 0.94 0.28 -324% -426% 0.96
naval_propulsion_plant 108% 1% 52% 0.96 1.03 0.30 -391% -469% 0.95
miami_housing -7% 0% -7% 0.95 1.22 0.15 -11% -20% 0.90
miami_housing -10% -1% -10% 0.95 1.24 0.15 -6% -17% 0.90
miami_housing -8% -1% -8% 0.95 1.25 0.15 -13% -23% 0.89
miami_housing -9% -1% -10% 0.95 1.22 0.15 -9% -19% 0.90
miami_housing -9% 0% -9% 0.94 1.24 0.15 -10% -22% 0.89
kin8nm -10% 1% -12% 0.95 2.75 0.54 -36% -40% 0.46
kin8nm -8% 0% -9% 0.94 2.66 0.52 -33% -36% 0.49
kin8nm -11% 3% -15% 0.95 2.76 0.54 -32% -35% 0.50
kin8nm -9% 2% -12% 0.94 2.67 0.52 -36% -40% 0.48
kin8nm -10% 1% -13% 0.95 2.73 0.53 -33% -36% 0.49
concrete_compressive_strength -18% -3% -18% 0.89 1.67 0.34 -78% -110% 0.74
concrete_compressive_strength -5% -3% -3% 0.89 1.51 0.30 -71% -97% 0.76
concrete_compressive_strength -6% -6% 0% 0.87 1.42 0.29 -63% -87% 0.77
concrete_compressive_strength -6% -3% -4% 0.90 1.47 0.30 -63% -91% 0.77
concrete_compressive_strength -6% 1% -7% 0.92 1.59 0.32 -61% -88% 0.77
cars -35% 5% -60% 0.95 1.01 0.17 6% 3% 0.95
cars -25% 1% -34% 0.91 0.83 0.14 8% 4% 0.95
cars -19% 0% -24% 0.92 0.80 0.14 4% -2% 0.95
cars -24% -3% -28% 0.90 0.86 0.15 2% 0% 0.95
cars -21% -3% -22% 0.91 0.91 0.15 -3% 1% 0.94
energy_efficiency 16% -8% 21% 0.90 0.18 0.05 0% 0% 1.00
energy_efficiency 6% -2% 6% 0.97 0.21 0.06 4% 1% 1.00
energy_efficiency 8% -4% 11% 0.95 0.21 0.06 5% 2% 1.00
energy_efficiency 3% -3% 5% 0.96 0.21 0.06 1% 0% 1.00
energy_efficiency 17% -4% 18% 0.95 0.19 0.06 5% 0% 1.00
california_housing -2% -2% 1% 0.95 2.38 0.57 -16% -22% 0.70
california_housing 3% -2% 5% 0.95 2.31 0.55 -13% -19% 0.72
california_housing 2% -2% 4% 0.95 2.29 0.55 -15% -20% 0.71
california_housing 0% -3% 3% 0.95 2.32 0.55 -16% -22% 0.71
california_housing 1% -3% 4% 0.95 2.32 0.55 -14% -20% 0.72
airfoil_self_noise -25% -5% -26% 0.94 2.17 0.46 -166% -184% 0.74
airfoil_self_noise 0% -1% 1% 0.98 1.78 0.38 -59% -80% 0.89
airfoil_self_noise -15% -1% -16% 0.98 1.98 0.42 -97% -106% 0.84
airfoil_self_noise -28% -5% -32% 0.93 2.20 0.47 -169% -213% 0.67
airfoil_self_noise -14% -3% -13% 0.96 1.93 0.41 -98% -125% 0.84
QSAR_fish_toxicity -16% -2% -17% 0.88 2.45 0.38 -6% -10% 0.53
QSAR_fish_toxicity -10% -2% -9% 0.89 2.43 0.38 -7% -12% 0.52
QSAR_fish_toxicity 3% -6% 9% 0.88 2.38 0.37 -8% -12% 0.50
QSAR_fish_toxicity -12% -4% -9% 0.88 2.41 0.38 -5% -10% 0.55
QSAR_fish_toxicity 8% -9% 16% 0.84 2.03 0.32 -6% -13% 0.53
@@
space_ga 5% -2% 7% 0.95 2.11 0.23 0% -1% 0.74
space_ga 3% -1% 3% 0.96 2.13 0.23 -1% -1% 0.75
space_ga 6% -1% 7% 0.94 2.04 0.22 -3% -3% 0.74
space_ga 5% -2% 6% 0.95 2.15 0.23 0% 0% 0.76
space_ga 10% 0% 9% 0.95 2.06 0.22 1% 0% 0.76
cpu_activity -18% 1% -22% 0.96 0.58 0.11 4% 2% 0.98
cpu_activity -3% 0% -2% 0.95 0.51 0.10 8% 7% 0.98
cpu_activity -9% 0% -10% 0.95 0.53 0.10 6% 5% 0.98
cpu_activity -6% 0% -7% 0.95 0.54 0.10 6% 5% 0.98
cpu_activity -2% 0% -2% 0.95 0.53 0.10 8% 5% 0.98
naval_propulsion_plant 32% 1% 24% 0.95 0.23 0.07 -44% -48% 1.00
naval_propulsion_plant 0% 1% -2% 0.96 0.31 0.09 -9% -11% 1.00
naval_propulsion_plant 7% 0% 5% 0.95 0.24 0.07 -35% -13% 1.00
naval_propulsion_plant -8% 1% -11% 0.96 0.28 0.08 -86% -63% 1.00
naval_propulsion_plant -13% 2% -16% 0.96 0.34 0.10 -51% -45% 0.99
miami_housing -38% 0% -60% 0.95 1.18 0.15 -6% 1% 0.91
miami_housing -37% 0% -59% 0.95 1.16 0.14 -2% 7% 0.91
miami_housing -43% 0% -74% 0.95 1.25 0.15 -4% 6% 0.91
miami_housing -38% 1% -62% 0.95 1.15 0.14 -2% 3% 0.92
miami_housing -33% -1% -47% 0.95 1.10 0.14 -1% 4% 0.91
kin8nm 10% 1% 9% 0.95 1.03 0.20 4% 5% 0.93
kin8nm 2% 1% 1% 0.96 1.07 0.21 10% 10% 0.94
kin8nm 4% 1% 3% 0.95 1.06 0.21 10% 6% 0.94
kin8nm 10% -3% 12% 0.93 0.99 0.19 6% 5% 0.93
kin8nm 15% -1% 13% 0.94 1.03 0.20 7% 7% 0.93
concrete_compressive_strength 7% -3% 9% 0.91 1.24 0.25 18% 19% 0.92
concrete_compressive_strength 18% 1% 14% 0.97 1.39 0.28 9% 12% 0.91
concrete_compressive_strength 2% 1% 1% 0.97 1.63 0.33 9% 15% 0.91
concrete_compressive_strength 7% -5% 11% 0.94 1.62 0.33 4% 10% 0.90
concrete_compressive_strength 19% -3% 18% 0.93 1.26 0.26 18% 20% 0.91
cars 5% -9% 13% 0.88 0.66 0.11 -5% -2% 0.95
cars 0% -3% 2% 0.93 0.80 0.14 -6% -2% 0.95
cars -9% 0% -10% 0.93 0.80 0.14 1% 1% 0.96
cars -8% -3% -4% 0.93 0.83 0.14 -2% -1% 0.95
cars 25% -1% 21% 0.94 0.79 0.13 -4% 5% 0.95
energy_efficiency 76% -3% 46% 0.94 0.16 0.04 31% 31% 1.00
energy_efficiency 46% 7% 25% 0.99 0.19 0.06 35% 32% 1.00
energy_efficiency 34% -1% 27% 0.95 0.17 0.05 28% 32% 1.00
energy_efficiency 29% -3% 25% 0.95 0.16 0.05 31% 32% 1.00
energy_efficiency 79% -3% 45% 0.96 0.17 0.05 37% 38% 1.00
california_housing -13% 1% -16% 0.95 1.88 0.45 8% 11% 0.82
california_housing -8% 0% -9% 0.95 1.78 0.42 7% 8% 0.81
california_housing -16% 1% -20% 0.95 1.83 0.44 7% 9% 0.82
california_housing -$-1$2% 0.94 1.82 0.43 7% 8% 0.81
california_housing -12% 0% -13% 0.94 1.76 0.42 7% 9% 0.82
airfoil_self_noise 98% -2% 51% 0.97 0.82 0.18 21% 19% 0.97
airfoil_self_noise 33% 0% 24% 0.98 0.90 0.19 17% 12% 0.97
airfoil_self_noise 36% -3% 28% 0.95 0.79 0.17 11% 9% 0.97
airfoil_self_noise 83% -4% 47% 0.95 0.76 0.16 22% 22% 0.97
airfoil_self_noise 93% -1% 49% 0.98 0.88 0.19 18% 18% 0.97
QSAR_fish_toxicity -11% 2% -16% 0.91 2.53 0.39 5% 11% 0.51
QSAR_fish_toxicity 3% 2% 1% 0.91 2.38 0.37 0% -1% 0.57
QSAR_fish_toxicity 7% -1% 7% 0.91 2.31 0.36 6% 11% 0.54
QSAR_fish_toxicity 18% 2% 14% 0.90 2.10 0.33 -4% 6% 0.52
QSAR_fish_toxicity 9% -1% 9% 0.83 1.84 0.29 10% 9% 0.61
@@
space_ga 4% -4% 7% 0.93 2.82 0.31 -11% -13% 0.35
space_ga 2% -1% 3% 0.94 2.97 0.32 -6% -6% 0.45
space_ga -4% -3% -2% 0.94 3.06 0.33 -12% -11% 0.44
space_ga -1% -1% -1% 0.95 2.97 0.32 -1% -6% 0.46
space_ga 2% 0% 2% 0.95 2.77 0.30 -5% -11% 0.54
cpu_activity -9% 1% -11% 0.95 0.89 0.17 -7% -23% 0.94
cpu_activity -26% 1% -44% 0.93 1.02 0.19 -1% -26% 0.81
cpu_activity -15% -2% -14% 0.89 0.80 0.15 -30% -30% 0.73
cpu_activity -1% -1% -3% 0.95 0.86 0.16 36% 5% 0.88
cpu_activity -19% 2% -26% 0.98 1.12 0.21 -2% -11% 0.81
naval_propulsion_plant -8% 0% -11% 0.98 2.93 0.86 -45% -90% 0.73
naval_propulsion_plant 31% -2% 24% 0.93 1.50 0.44 5% -6% 0.84
naval_propulsion_plant 18% -2% 17% 0.96 2.06 0.61 -30% -42% 0.83
naval_propulsion_plant 53% -5% 37% 0.87 1.27 0.37 6% 6% 0.76
naval_propulsion_plant -5% -3% -6% 0.93 2.22 0.65 -7% -11% 0.70
miami_housing -10% -2% -11% 0.94 1.45 0.18 -7% -12% 0.80
miami_housing -38% 0% -63% 0.93 1.86 0.23 -35% -35% 0.54
miami_housing -26% -1% -33% 0.95 1.74 0.22 -10% -19% 0.76
miami_housing 1% -3% 2% 0.95 1.35 0.17 -14% -20% 0.79
miami_housing -17% 1% -22% 0.94 1.64 0.20 -4% -4% 0.72
kin8nm -6% 0% -8% 0.97 3.36 0.66 0% 1% 0.40
kin8nm -15% 1% -22% 0.97 3.46 0.68 -6% -8% 0.39
kin8nm -8% 3% -12% 0.95 3.00 0.59 1% 2% 0.41
kin8nm -9% 2% -14% 0.96 3.11 0.61 -5% -7% 0.44
kin8nm -8% 1% -9% 0.95 3.23 0.63 -5% -6% 0.37
concrete_compressive_strength -8% -3% -8% 0.95 2.81 0.57 -19% -22% 0.48
concrete_compressive_strength 7% -5% 11% 0.93 2.79 0.56 -37% -44% 0.57
concrete_compressive_strength 14% -3% 13% 0.97 2.69 0.55 -4% -9% 0.64
concrete_compressive_strength 4% -1% 3% 0.95 2.83 0.57 -15% -24% 0.50
concrete_compressive_strength -11% 0% -14% 0.97 3.01 0.61 -4% -3% 0.69
cars 31% -3% 24% 0.93 1.91 0.32 -9% -16% 0.75
cars 51% -7% 38% 0.93 1.84 0.31 -37% -25% 0.81
cars -16% -11% -17% 0.89 2.17 0.37 -68% -80% 0.55
cars -13% -3% -21% 0.94 2.18 0.37 -8% -14% 0.76
cars -5% -12% 5% 0.83 2.00 0.34 -28% -39% 0.40
energy_efficiency -11% 0% -13% 1.00 1.15 0.33 -15% -24% 0.97
energy_efficiency 20% -10% 25% 0.89 0.84 0.24 -705% -520% 0.76
energy_efficiency 42% -1% 23% 0.95 0.78 0.23 -37% -39% 0.94
energy_efficiency 61% -4% 40% 0.94 0.70 0.20 -16% -44% 0.96
energy_efficiency 20% -3% 16% 0.97 1.12 0.33 -155% -137% 0.93
california_housing -14% 0% -17% 0.96 2.67 0.64 -3% -3% 0.65
california_housing 0% -2% 0% 0.96 2.41 0.57 -11% -13% 0.73
california_housing -7% -1% -8% 0.96 2.48 0.59 0% 1% 0.73
california_housing 2% -3% 3% 0.92 2.28 0.54 -4% -7% 0.61
california_housing -12% -1% -16% 0.97 2.63 0.63 -1% -3% 0.73
airfoil_self_noise -30% -2% -43% 0.97 3.95 0.84 -94% -110% 0.50
airfoil_self_noise -18% -1% -22% 0.91 2.98 0.64 -19% -29% 0.38
airfoil_self_noise -23% -4% -25% 0.91 3.49 0.75 -43% -48% 0.19
airfoil_self_noise -31% 1% -49% 0.97 3.61 0.77 -14% -8% 0.52
airfoil_self_noise -17% -1% -21% 0.97 3.81 0.81 -14% -22% 0.43
QSAR_fish_toxicity -8% 3% -14% 0.98 4.22 0.66 4% 4% 0.43
QSAR_fish_toxicity -6% 1% -7% 0.97 3.93 0.61 -9% -3% 0.35
QSAR_fish_toxicity 8% -5% 10% 0.88 2.59 0.40 -4% -5% 0.38
QSAR_fish_toxicity -5% -5% -5% 0.88 2.83 0.44 -7% -3% 0.35
QSAR_fish_toxicity 12% -11% 15% 0.85 2.58 0.40 -5% -3% 0.30